If 32^(a+b) = 16^(a+2b), then a =
A) b
B) 2b
C) 3b
D) b + 2
E) b - 2
I thought you're supposed to try to get a common base then try to solve but I can't seem to get it..
Thanks!

Answer by jim_thompson5910(35256)

32^(a+b) = 16^(a+2b)

(2^5)^(a+b) = (2^4)^(a+2b)

2^(5(a+b)) = 2^(4(a+2b))

5(a+b) = 4(a+2b)

5a+5b = 4a+8b

5a-4a = 8b-5b

a = 3b

So the answer is choice C